# Leibniz rule for wedge product in discrete exterior calculus on general polygonal meshes

### v1

@@ 6 Conclusion
We have presented a new definition of discrete wedge product on polygonal 2–pseudomanifolds (Definition 15) and proved that this polygonal wedge product satisfies Leibniz product rule with discrete exterior derivative (Lemma 3). We have also shown that the polygonal wedge product is a cup product (Theorem 4). Moreover, in Proposition 5 we have proved that the polygonal wedge product shares further properties with the wedge product of differential forms, such skew–commutativity.
Declarations
Funding The author was supported in part by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq): processos 140396/2013-6, 300735/2017-0. The author was supported also in part by PRIMUS/25/SCI/022.
Data Availability Not applicable.
Conflict of interest The author declares no competing interests.

### v3

@@ 6 Conclusion
We have presented a new definition of discrete wedge product on polygonal 2–pseudomanifolds (Definition 5.4) and proved that this polygonal wedge product satisfies the Leibniz product rule with discrete exterior derivative (Lemma 5.1). We have also shown that the polygonal wedge product is a cup product (Theorem 5.1). Moreover, in Proposition 5.3 we have proved that the polygonal wedge product shares further properties with the wedge product of differential forms, such skew–commutativity.
Declaration of competing interest
The author declares that she has no competing interests.
Acknowledgement
The author was supported in part by Conselho Nacional de Desenvolvimento Científico e Tecnológico: processos 140396/2013-6, 300735/2017-0; and in part by the Czech Science Foundation Grant 25-15714S.